OpenStudy (anonymous):

How do you get the slope of the line passing through points (1,3) and (-3, -5)?

OpenStudy (tkhunny):

You calculate it! \(\dfrac{Rise}{Run}=\dfrac{\Delta Y}{\Delta X}=\dfrac{Change\;in\;Y}{Change\;in\;X}\). Let's see what you get.